wiki:Research/Instrumentation/Meetings/2019/07-10

SmartPhantom meeting: Wednesday 10th July 2019, 11:00 BST

Meeting details

Teleconference

Phone details:

UK: 08082 380274
Austria: 0800 006612 

PIN: 5904750902

Full list of telephone contacts may be found here.

Notes:

Present: JB, AK, KL, LC

  1. Notes and actions from last meeting:
    • AK: Complete order for clear fibre.
      • In hand: Purchase order raised. Saint Gobain has acknowledged receipt. Order is expected in 6 weeks.
    • LC: Re-issue drawings 'for the record' when ready.
      • Done.
    • HTL: With KL, review simulation of doublet layers potted in Araldite.
      • Done. KL discussed with HTL. Small differences in simulation were acceptable. Adopt potting in black epoxy as baseline.
    • HTL: With KL, review estimate of light yield from simulation.
      • Done. HTL had revised light-yield calculation and discussed with KL. Now seems OK for present.
    • HTL: Inject 90Sr electrons to estimate test-rig light yeild.
      • In progress. HTL has simulation that gives light yield and has found 90Sr spectrum. Work in progress.
    • KL: Liaise to develop concept for polishing and light-exclusion from light guides.
      • Done for light exclusion; LC has a concept that is discussed below.
      • Stands: Polishing specification. AK mentioend that Saint Gobain specification may include spec on polishing. Need to make a decision on strategy.
    • LC: Re-evaluate use of PTFE based on discussion with R12.
      • Done. See below.
    • KL: Work within Imperial to find assembly space.
      • On going. Could be 527.
    • KL: Seek meeting with S.Gruber to get her input into the design.
      • Stands.
  1. Design update: LC
  • Points made in LC update were discussed. We noted:
    • That the chemical composition of the black dye (and the epoxy) was required. LC will discuss with S.Camfer of the Resin Lab at RAL.
    • Need a mechanical design for the potting of the fibres emerging from the optical connector and for the connector that mates with the camera. LC agreed to look at these.
    • There may be a requirement for a tensioner, particularly if we move to production of a large number of detectors.
  1. Update on progress of preparations for manufacture: LC & discussion
    • 6 quotes had been obtained for manufacture of aluminium parts. The cheepest was £1069 (+ VAT?). LC will check with JT to understand whether the company is acceptable. If it is, we agreed to place the order.
    • KL will email FW to work out how best to place the order.
    • We noted that winding most probably would take place at Imperial.
  1. Simulation update: HTL
    • No update as HTL on shift at MedAustron.
  1. Optical readout update: JB
    • Status report 03Jul19
    • We agreed to go with 'option 1', the commercial unit. JB will move to order of parts. AK will send relevant project code to JB.
  1. Next steps
    Move to procurement of components so that we can begin to work towards manufacture of first prototype.
  1. DONM
    • 24Jul19;
  1. AoB
    We agreed to treat the presentation to SG as a mini review of the project so far.

Summary of actions:

  • HTL: Inject 90Sr electrons to estimate test-rig light yeild.
  • KL: Liaise to develop concept for polishing.
  • KL: Work within Imperial to find assembly space.
  • KL: Seek meeting with S.Gruber to get her input into the design.
  • LC: Obtain chemical composition of resin and black dye.
  • LC: Develop concept for optical connector at camera end of lightguide and for potting of light guide at detector end.
  • KL: Liaise with FW to work out how to place aluminium-part order.
  • JB: Purchase parts for optical readout system.
  • AK: Provide JB with project code for purchase of optical readout components.


Agenda:

  1. Notes and actions from last meeting:
  • AK: Complete order for clear fibre.
  • LC: Re-issue drawings 'for the record' when ready.
  • HTL: With KL, review simulation of doublet layers potted in Araldite.
  • HTL: With KL, review estimate of light yield from simulation.
  • HTL: Inject 90Sr electrons to estimate test-rig light yeild.
  • KL: Liaise to develop concept for polishing and light-exclusion from light guides.
  • LC: Re-evaluate use of PTFE based on discussion with R12.
  • KL: Work within Imperial to find assembly space.
  • KL: Seek meeting with S.Gruber to get her input into the design.
  1. Design update: LC
  1. Update on progress of preparations for manufacture: LC & discussion
  1. Simulation update: HTL
  1. Optical readout update: JB
  1. Next steps
  1. DONM
  1. AoB

Last modified 2 years ago Last modified on Jul 19, 2019 7:44:19 PM

Attachments (6)

Download all attachments as: .zip